OpenStudy (anonymous):

ΔCAR has coordinates C (2, 4), A (1, 1), and R (3, 0). A translation maps point C to C' (3, 2). Find the coordinates of A' and R' under this translation. A' (4, -2); R' (2, -1) A' (-2, 2); R' (2, -2) A' (2, -1); R' (4, -2) A' (-1, 0); R' (-2, 2)

OpenStudy (anonymous):

If this is translation, that means that the triangle's only been moved a certain distance in each direction, right? Find the difference between C and C', and apply it to each point. \[T=\left( \Delta x,\Delta y \right) = \left( 1,-2 \right)\]\[A'=A+T\]\[R'=R+T\]